A regulation basketball has a 50 cm diameter
and may be approximated as a thin spherical
shell.
How long will it take a basketball starting
from rest to roll without slipping 1.9 m down
an incline that makes an angle of 77.0
◦ with
the horizontal? The acceleration of gravity is
9.81 m/s
2
.
Answer in units of s.
1
Expert's answer
2020-02-03T13:59:37-0500

The moment of inertia of the ball is


"I=\frac{2}{5}mr^2"

The initial height of the ball is


"h=1.9\sin{77}"

The initial potential energy is


"mgh=mg(1.9\sin{77})"

From the conservation of energy:


"mgh=0.5mv^2 +0.5I\u03c9^2 \u2192mgh=0.5mv^2 +0.5(0.4mr^2 )"

"gh=0.5v ^\n2\n +0.5(0.4v^\n2\n )=0.7v^ \n2"

"9.81(1.9\sin{77})=0.7v^2"

"v=5.094\frac{m}{s}"

"t=\frac{2s}{v}=\frac{2(1.9)}{5.094}=0.746\ s"